# SeniorTalk: A Chinese Conversation Dataset with Rich Annotations for Super-Aged Seniors

### v1

@@ 4.2.3 Results and Analysis
Second, our empirical analysis on speaker diarization reveals that ECAPA-TDNN significantly outperforms both ResNet-TDNN and x-vector-based systems. However, Table 3 demonstrates that these performance gaps diminish in speaker verification tasks, with ResNet-TDNN even surpassing ECAPA-TDNN under PLDA scoring. We attribute this discrepancy to two primary factors: the shift to a conversational dataset and ECAPA-TDNN’s inherent architectural advantages in noise resilience. While ECAPA-TDNN is typically trained on clean speech corpora, our experiments evaluate its robustness in realistic scenarios characterized by background noise and overlapping speech. The model’s noise-resilient components enhance its feature extraction capabilities under such adverse conditions, which proves critical in diarization tasks but less impactful in controlled speaker verification settings. This contrast highlights the importance of conversational datasets in assessing model robustness, as they better reflect real-world challenges that differentiate model performance.
Third, our experiments reveal notably high DER and confusion errors. Our analysis attributes this phenomenon to two dataset-specific factors: (1) a pronounced gender imbalance (1:3 female-to-male ratio) contrasting standard benchmarks’ balanced distributions, and (2) age-related vocal changes in elderly speakers that reduce the saliency of secondary sexual voice characteristics. These combined effects create systematic challenges for speaker identity separation, particularly in conversational contexts where demographic diversity and physiological aging patterns naturally occur.
4.3 Speech Recogition
@@
Table 6: Character Error Rate (CER) (%) of supervised pre-trained baselines in both zero-shot and fine-tuning settings.
4.3.1 Metrics
$CER=\frac{S+D+I}{N},$ (2)
4.3.2 Baselines
We employ the open-source wenet toolkit (Yao et al., 2021) as our training framework, selecting Transformer, Conformer, and E-Branchformer as our baselines. All models are trained using a combined Connectionist Temporal Classification (CTC) and Attention-based Encoder-Decoder (AED) approach. We fine-tune the hyperparameters of these three models to ensure comparable parameter counts while achieving optimal performance. Detailed hyperparameter configurations can be found in the Appendix B.2.1 and B.2.2.
  • Conformer: The Conformer (Gulati et al., 2020) model integrates convolutions with self-attention for ASR, sandwiched between two feed-forward layers.
  • E-Branchformer: Proposed by Kwangyoun Kim et al. (Kim et al., 2022), E-Branchformer builds upon the Branchformer (Peng et al., 2022), which attains performance levels comparable to Conformer. E-Branchformer improves on this framework by implementing a novel merging strategy and integrating additional point-wise modules.
  • Paraformer: Proposed by Gao et al. (Gao et al., 2022), Paraformer is a fast and accurate parallel transformer model that leverages a continuous integrate-and-fire (CIF) (Dong and Xu, 2020) predictor to estimate the number of tokens and generate hidden representations. This pre-trained model is trained on a non-public, industry-grade dataset comprising 60,000 hours of Chinese ASR data.
  • Whisper: Whisper (Radford et al., 2023) is a Transformer-based multilingual ASR model developed by OpenAI, trained on 68,000 hours of labeled speech data. We examine various versions of Whisper, ranging from tiny to large-v2, with model sizes varying from 39M to 1.55Bhttps://github.com/openai/whisper.
4.3.3 Results and Analysis
Models Trained from Scratch
Accent Intensity Impact: As shown in Table 5, CER increases with accent intensity: Moderate Accent yields higher CER than Light Accent, which in turn exceeds No Accent. This trend highlights the growing recognition challenge as accents become more pronounced.
Regional Variations: The E-Branchformer achieves a 0.97% lower CER in the South than in the North, with the Conformer and Transformer showing improvements of 0.19% and 1.74%, respectively. These differences indicate slightly higher recognition difficulty in the North, although overall performance remains comparable. At the provincial level, substantial variations in CER are observed across different provinces. For detailed analysis, please refer to the AppendixB.3.
@@ 4.3.3 Results and Analysis
Table 6 presents comparative CER results for two model families: 1) our Wenet-finetuned Paraformer-large architecture (Yao et al., 2021), and 2) a series of Whisper models (Radford et al., 2023) adapted through whisper-flamingo fine-tuninghttps://github.com/roudimit/whisper-flamingo. The analysis reveals two key findings:
First, although Whisper models show modest CER improvements with larger parameter counts (except for small), all variants have CERs exceeding 50% in zero-shot senario. This performance gap suggests potential domain mismatch between Whisper’s training data distribution and elderly speech characteristics, possibly due to underrepresentation of senior voices in Whisper’s pretraining corpus. Such mismatch may explain the frequently observed hallucination patterns in elderly speech recognition results. Notably, our targeted fine-tuning reduces CERs substantially, improving cer of Whisper-large-v3 from 57.74 to 23.84%, demonstrating both the challenge level of our elderly speech dataset and its utility for domain adaptation.
Second, the Paraformer-large model achieves 14.91% CER in zero-shot evaluation and 14.41% CER after fine-tuning , outperforming all Whisper variants by significant margins. This advantage likely stems from Paraformer’s pretraining on 60,000 hours of proprietary Chinese speech data encompassing diverse regional accents, age groups, and speaking styles. This demonstrates that Paraformer has better generalization ability in the field of Chinese speech recognition.
@@
Table 7: Evaluation results of speech editing models trained from scratch.
@@ 4.4 Speech Editing
4.4.1 Baselines
We employ the Speech Editing Toolkit https://github.com/pyannote/pyannote-audio framework to implement our models, which include CampNet(Wang et al., 2022), EditSpeech(Tan et al., 2021), A3T(Bai et al., 2022) , and FluentSpeech(Jiang et al., 2023). The detailed hyperparams are described in Appendix B.2.4.
4.4.2 Metrics
We utilize Mel-Cepstral Distortion (MCD) Kubichek (1993), Short-Time Objective Intelligibility (STOI) Taal et al. (2010), and Perceptual Evaluation of Speech Quality (PESQ) Rix et al. (2001), which are commonly used objective metrics for assessing speech generation quality.
@@ 4.4.3 Performance Analysis
The results presented in Table 7 indicate that FluentSpeech achieved the best performance, with all three objective metrics falling within acceptable ranges. In relevant research (Jiang et al., 2023; Bai et al., 2022; Wang et al., 2022; Liu et al., 2024), these experimental values are widely recognized as acceptable. This suggests that our dataset is suitable for generative tasks.
@@ 5 Conclusion
Limitations
While SeniorTalk advances research on elderly speech processing, several limitations should be noted. First, while our dataset includes 202 participants across 16 provinces, its scale remains modest compared to the linguistic diversity and population size of China. Expanding the dataset to include thousands of speakers would better capture regional accents, dialects, and individual vocal variability. Second, the current age range could be extended to include more super-elderly individuals (e.g. 85+ or centenarians), as advanced aging often correlates with unique vocal patterns due to physiological decline, cognitive changes, or age-related pathologies. Third, geographic coverage could be broadened to encompass underrepresented regions (e.g., rural areas or ethnic minority regions) to mitigate potential demographic biases. Additionally, longitudinal data capturing vocal aging trajectories could further enhance the dataset’s utility. These extensions would strengthen the generalizability of voice technologies for aging populations and address the complex interplay of age, health, and regional linguistic diversity.
Ethics Statement
This framework ensures compliance with international research ethics standards (Declaration of Helsinki, GDPR principles) while balancing scientific utility with participant protection in vulnerable populations.
@@ A.2 Topic
A.3 Annotation
@@ 14 Hunan 47 Male Bachelor
Table 10: Annotator Information
A.3.1 Annotator
@@ B.1 Data Split
Encoder Accum Grad Batch size LR Warmup Epochs
Transformer 4 4 1.00E-03 5000 100
Conformer 4 4 1.00E-03 5000 60
E-Branchformer 4 4 5.00E-04 5000 60
Table 12: Hyperparameters for training ASR models from scratch.
Model Accum Grad Batch size Learning rate Warmup Training steps
Paraformer-large 4 28 5.00E-04 5000 101579
Whisper-large 1 8 5.00E-6 1000 90000
Whisper-medium 1 3 6.25e-6 1000 225000
Whisper-small 1 4 1.25e-5 1000 225000
Table 13: Hyperparameters for fine-tuning pre-trained ASR models.
Model Batch size LR Schedule Init LR Base LR Epochs
ECAPA-TDNN 128 Cyclic 5.00E-03 1.00E-08 20
ResNet-TDNN 128 Cyclic 5.00E-03 1.00E-08 20
x-vector 128 Linear 5.00E-03 1.00E-04 20
Table 14: Hyperparameters for training speaker verification models.
Model Params Batch size LR Warmup Training steps
CampNet 21.22M 16 2.00E-04 8000 200000
EditSpeech 48.15M 16 2.00E-04 8000 200000
A3T 14.86M 16 2.00E-04 8000 200000
FluentSpeech 23.86M 30 2.00E-04 8000 200000
Table 15: Hyperparameters for training speaker editing models.
Model Shanghai Guangxi Sichuan Anhui Zhejiang Jiangsu Hunan
Paraformer-fintuned 13.68 4.88 9.62 16.11 12.42 16.57 14.6
Paraformer 13.26 6.24 10.07 17.91 12.3 16.61 13.18
Brachformer 27.37 10.82 25.94 36.58 29.83 36.68 39.94
Conformer 29.11 11.91 27.11 38.01 30.17 39.44 40.65
Tranformer 42.27 26.91 41.78 52.3 46.53 51.82 54.9
Table 16: Model Performance in Southern Provinces
Model Beijing Heilongjiang Liaoning Hebei Henan
Paraformer-fintuned 7.38 5.85 27.08 7.71 20.93
Paraformer 6.8 5.38 23.85 9.41 24.19
Brachformer 23.65 21.21 55.31 19.63 40.85
Conformer 25.15 20.27 56.88 20.44 41.59
Tranformer 43.01 39.61 67.58 36.61 56.03
Table 17: Model Performance in Northern Provinces
@@ B.1.1 Speaker Verification and Diarization
For the speaker verification experiments, we first segment the original dialogue dataset into sentence-level units. Given that the dataset includes precise timestamps for each sentence and specific audio markers for events such as overlapping speech and noise, we extract clean audio segments devoid of these special sound events and speaker overlaps. From the 101 dialogues in the dataset, we randomly select 10 dialogues, using the corresponding segmented audio at the sentence level as our test dataset. The remaining audio is split, with 10% randomly allocated for the validation set and 90% designated for the training set. The detailed information of the datasplit is shown in 11. After the datasplit, we create 20,000 carefully balanced verification pairs (50% genuine vs. 50% impostor pairs) from the test set. This trial composition ensures uniform coverage of both intra-speaker ${(spk_{i},spk_{j})}$ and inter-speaker pairs ${(spk_{i},spk_{j})}$. The detailed datasplit information is shown in table 11. For speaker diarization, the dataset split is identical to that of the speaker verification task, except that we use dialogues instead of sentences.

### v2

@@ 4.2.3 Results and Analysis
Second, as shown in Table 4, ECAPA-TDNN achieves significantly superior performance over both ResNet-TDNN and x-vector-based baselines on the speaker diarization task. Moreover, consistent with these findings, the ECAPA-TDNN demonstrates consistently lower EER across both PLDA and cosine similarity senarios on the speaker verification task as shown in table 3. These results highlight the ECAPA-TDNN’s robustness, confirming its strong generalization ability not only in clean, utterance-level speaker verification tasks but also in more challenging, real-world conversational speaker diarization scenarios involving increased noise and overlapping speech.
Third, our experiments reveal notably high DER and confusion errors. Our analysis attributes this phenomenon to two dataset-specific factors: (1) a pronounced gender imbalance (1:2 female-to-male ratio) contrasting standard benchmarks’ balanced distributions, and (2) age-related vocal changes in elderly speakers that reduce the saliency of secondary sexual voice characteristics. These combined effects create systematic challenges for speaker identity separation, particularly in conversational contexts where demographic diversity and physiological aging patterns naturally occur.
@@
4.3 Speech Recogition
4.3.1 Metrics
$CER=\frac{S+D+I}{N},$ (2)
4.3.2 Baselines
We employ the open-source Wenet toolkit (wenet) as our training framework, selecting Transformer, Conformer, and E-Branchformer as our baselines. All models are trained using a combined Connectionist Temporal Classification (CTC) and Attention-based Encoder-Decoder (AED) approach. We fine-tune the hyperparameters of these three models to ensure comparable parameter counts while achieving optimal performance. Detailed hyperparameter configurations can be found in the Appendix B.2.1 and B.2.2.
  • Conformer: The Conformer (conformer) model integrates convolutions with self-attention for ASR, sandwiched between two feed-forward layers.
  • E-Branchformer: Proposed by Kwangyoun Kim et al. (E-Branchformer), E-Branchformer builds upon the Branchformer (Branchformer), which attains performance levels comparable to Conformer. E-Branchformer improves on this framework by implementing a novel merging strategy and integrating additional point-wise modules.
  • Paraformer: Proposed by Gao et al. (paraformer), Paraformer is a fast and accurate parallel transformer model that leverages a continuous integrate-and-fire (CIF) (cif) predictor to estimate the number of tokens and generate hidden representations. This pre-trained model is trained on a non-public, industry-grade dataset comprising 60,000 hours of Chinese ASR data.
  • Whisper: Whisper (whisper)  [^9] https://github.com/openai/whisper is a Transformer-based multilingual ASR model developed by OpenAI, trained on 680,000 hours of labeled speech data. We examine various versions of Whisper, ranging from tiny to large-v2, with model sizes varying from 39M to 1.55B.
4.3.3 Results and Analysis
Models Trained from Scratch
Accent Intensity Impact: As shown in Table 5, CER increases with accent intensity: Moderate Accent yields higher CER than Light Accent, which in turn exceeds No Accent. This trend highlights the growing recognition challenge as accents become more pronounced. The duration distribution of utterances across different accent levels is presented in Appendix A.3.
Regional Variations: The E-Branchformer achieves a 0.97% lower CER in the South than in the North, with the Conformer and Transformer showing improvements of 0.19% and 1.74%, respectively. These differences indicate slightly higher recognition difficulty in the North, although overall performance remains comparable. At the provincial level, substantial variations in CER are observed across different provinces. For detailed analysis, please refer to the AppendixB.3.1.
Table 6: Character Error Rate (CER) (%) of the pretrained Paraformer-large model, along with various sizes of Whisper models (tiny, base, small, medium, and large-v3) under both zero-shot and fine-tuning settings.
@@ 4.3.3 Results and Analysis
Table 6 presents comparative CER results for two model families: 1) our Wenet-finetuned Paraformer-large architecture (wenet), and 2) a series of Whisper models (whisper) adapted through whisper-flamingo fine-tuning. [^10] https://github.com/roudimit/whisper-flamingo The analysis reveals two key findings:
First, although Whisper models show modest CER improvements with larger parameter counts (except for small), all variants have CERs exceeding 50% in zero-shot senario. This performance gap suggests potential domain mismatch between Whisper’s training data distribution and elderly speech characteristics, possibly due to underrepresentation of senior voices in Whisper’s pretraining corpus. Such mismatch may explain the frequently observed hallucination patterns in elderly speech recognition results. Notably, our targeted fine-tuning reduces CERs substantially, improving cer of Whisper-large-v3 from 57.74% to 23.84%, demonstrating both the challenge level of our elderly speech dataset and its utility for domain adaptation.
Second, the Paraformer-large model achieves 14.91% CER in zero-shot evaluation and 14.41% CER after fine-tuning , outperforming all Whisper variants by significant margins. This advantage likely stems from Paraformer’s pretraining on 60,000 hours of proprietary Chinese speech data encompassing diverse regional accents, age groups, and speaking styles. This demonstrates that Paraformer has better generalization ability in the field of Chinese elderly speech recognition.
Table 7: Objective Evaluation results of speech editing models trained from scratch.
@@
Table 8: Subjective Evaluation results of speech editing models trained from scratch.
Method Boundary-MOS MOS
CampNet 1.25 ± 0.33 1.17 ± 0.27
EditSpeech 2.42 ± 0.98 1.82 ± 0.91
A3T 3.00 ± 0.89 2.04 ± 0.74
FluentSpeech 4.72 ± 0.44 4.53 ± 0.57
@@ 4.4 Speech Editing
4.4.1 Baselines and Metrics
We employ the Speech Editing Toolkit  [^11] https://github.com/Zain-Jiang/Speech-Editing-Toolkit framework to implement our models, which include CampNet(campnet), EditSpeech(tan2021editspeech), A3T(a3t) , and FluentSpeech(flu). The detailed hyperparams are described in Appendix B.2.4.
For objective evaluation metrics, we utilize Mel-Cepstral Distortion (MCD) mcd, Short-Time Objective Intelligibility (STOI) stoi, and Perceptual Evaluation of Speech Quality (PESQ) pesq, which are commonly used objective metrics for assessing speech generation quality.
For subjective evaluation metrics, we conducted comprehensive human perceptual evaluation to assess speech editing quality. We randomly selected 50 audio samples and recruited three evaluators to rate each sample on a 1-5 scale using two metrics:
  • MOS (Mean Opinion Score): Evaluates naturalness of edited speech segments
  • Boundary-MOS: Our specialized metric measuring transition smoothness at editing boundaries
@@ 4.4.2 Performance Analysis
The results presented in Table 7 and 8 indicate that FluentSpeech achieved the best performance, with all objective metrics and subjective metrics falling within acceptable ranges. In relevant research (flu; a3t; campnet; liu2023fluenteditor), these experimental values are widely recognized as acceptable. This suggests that our dataset is suitable for generative tasks.
5 Limitations and Ethical Considerations
While SeniorTalk provides important data for elderly speech processing, it has notable limitations. The dataset includes 202 participants from 16 provinces but displays biases in gender, regional, and accent distributions. Additionally, the participant age range does not fully cover the super-elderly population (85+ years), and the overall dataset size remains relatively modest, which may restrict generalizability across diverse vocal characteristics. From an ethical perspective, rigorous safeguards were implemented: informed consent was obtained, recordings were conducted in controlled environments, and all personally identifiable information was anonymized. Participants received fair compensation in line with local economic conditions. To mitigate risks associated with misuse such as malicious voice synthesis, data access is limited to verified academic researchers under a non-commercial CC BY-NC-SA 4.0 license.
@@ 6 Conclusion
Acknowledgement
This work has been supported by the National Key R&D Program of China (Grant No.2022ZD0116307) and NSFChina (Grant No.62271270).
@@ Appendix A Supplementary Dataset Information
Table 9: Accent Intensity Annotation Guidelines
Level Annotation Criteria
No Accent (0) Phonetics: Fully standard pronunciation with clear articulation
Comprehensibility: Effortlessly understood by non-native listeners
Regional Features: No detectable regional phonological characteristics
Light Accent (1) Phonetics: Occasional non-standard vowels/consonants (<20% utterances)
Comprehensibility: Minor listening effort required for full understanding
Regional Features: Subtle but identifiable regional speech patterns
Moderate Accent (2) Phonetics: Frequent non-standard prosody/lexical stress (20-50% utterances)
Comprehensibility: Requires focused attention, occasional repetition needed
Regional Features: Strong regional phonological markers affecting intelligibility
Heavy Accent (3) Phonetics: Pervasive non-standard articulation (>50% utterances)
Comprehensibility: Frequent breakdowns requiring contextual guessing
Regional Features: Severely divergent from standard phonological norms
Table 10: Annotation Levels and Their Associated Tasks for Dataset Analysis
Annotation Level Annotation Dimension Associated Tasks Representative Instances
Speaker Metadata Demographic Age Elderly Speech Analysis 75
Geographic Origin (Province) Jiangsu, Henan
ID Card Gender Female/Male
Session Temporal Segmentation Speaker Diarization [48.475 - 73.582] spk_001
Overlapping Speech Speech Separation trans1(trans2)[+]
Utterance Raw Transcription Speech Recognition [Mandarin utterance]
Accent Intensity (0-3) Ordinal Classification Neutral (0) / Strong (3)
Token Special Markers Paralinguistic Analysis [MUSIC],[NOISE],[LAUGHTER]
@@ A.2 Topic
Figure 3: This distribution of topics.
A.3 Dataset Accent Distribution Analysis
This section presents the distribution characteristics of accent severity levels within our dataset. The accent annotation criteria follow the standardized framework detailed in table 9. As illustrated in Table 12, the distribution exhibits distinct patterns across different severity levels:
  • Light accent samples constitute the largest temporal proportion (110,419s) and account for 42,790 samples.
  • Moderate accent demonstrates the second-largest distribution with 17,782s duration and 8,226 samples.
  • Samples with no accent (5,951s/8,056) and heavy accent (370s/142) show minor representation.
The analysis reveals a pronounced concentration of data in the light accent category, which comprises over half of the total samples, followed by moderate, no, and heavy accents.
A.4 Annotation
@@ 14 Hunan 47 Male Bachelor
Table 12: Accent distribution
Accent Total Duration (seconds) Number of Samples
No 5,951 8,056
Light 110,419 42,790
Moderate 17,782 8,226
Heavy 370 142
A.4.1 Annotator
@@ A.4.3 Annotation Information
A.5 Licenses for assets
This work adheres to strict licensing compliance for all third-party assets. Below are the licensing terms for codebases and datasets utilized in our speech dataset creation paper:
  • WeNet: The WeNet framework [wenet] for end-to-end speech recognition is licensed under Apache 2.0, permitting modification and commercial use with attribution.
  • Whisper Flamingo: The Whisper Flamingo framework [whisper] for Whisper adaptation operates under BSD 3-Clause License, requiring copyright retention in derivative works.
  • SpeechBrain: SpeechBrain toolkit [speechbrain] utilizes Apache 2.0.
  • PyAnnote: PyAnnote toolkit[Bredin23, Plaquet23] utilizes MIT License which allows users to freely use, modify, and distribute the software, provided that the copyright notice and permission notice are retained.
  • Speech-Editing-Toolkit: The Speech-Editing-Toolkit [flu] includes no explicit software license but enforces strict usage prohibitions through appended declaration:
    “Any synthesis of recognizable voices (e.g., politicians, celebrities) without explicit consent violates ethical standards and copyright laws. Violators assume full legal liability.”
  • VoxCeleb Dataset: Training data for speaker modeling is governed by CC BY 4.0 [nagrani17_interspeech], mandating proper attribution for academic/commercial redistribution.
For reproducibility and open science, we release all components of our SeniorTalk under the following licenses:
  • SeniorTalk Dataset: Our conversational speech dataset is released under CC BY-NC-SA 4.0, restricting commercial applications while requiring share-alike terms for derivatives.
  • Codebase: Core infrastructure follows Apache License 2.0. Embedded third-party code retains original licensing.
@@ B.1 Data Split
Split # Spk. # Utt. Dur. (hrs) Avg. (s)
Train 0-182 48,591 30.47 2.26
Dev 0-182 5,398 3.40 2.27
Test 182-202 6,040 3.95 2.35
Sum 0-202 60,029 37.82 2.27
@@ B.1.1 Speaker Verification and Diarization
For the speaker verification experiments, we first segment the original dialogue dataset into sentence-level units. Given that the dataset includes precise timestamps for each sentence and specific audio markers for events such as overlapping speech and noise, we extract clean audio segments devoid of these special sound events and speaker overlaps. From the 101 dialogues in the dataset, we randomly select 10 dialogues, using the corresponding segmented audio at the sentence level as our test dataset. The remaining audio is split, with 10% randomly allocated for the validation set and 90% designated for the training set. The detailed information of the datasplit is shown in table 14. After the datasplit, we create 20,000 carefully balanced verification pairs (50% genuine vs. 50% impostor pairs) from the test set. This trial composition ensures uniform coverage of both intra-speaker $(spk_{i},spk_{j})$} and inter-speaker pairs $(spk_{i},spk_{j})$}. The detailed datasplit information is shown in table 14. For speaker diarization, the dataset split is identical to that of the speaker verification task, except that we use dialogues instead of sentences.
@@ B.2 Hyperparams
Table 17: Hyperparameters for training ASR models from scratch.
Encoder Accum Grad Batch size LR Warmup Epochs
Transformer 4 4 1e-3 5,000 100
Conformer 4 4 1e-3 5,000 60
E-Branchformer 4 4 5e-4 5,000 60
Table 18: Hyperparameters for fine-tuning pre-trained ASR models.
Model Accum Grad Batch size Learning rate Warmup Training steps
Paraformer-large 4 28 5e-4 5,000 101,579
Whisper-large 1 8 5e-6 1,000 90,000
Whisper-medium 1 3 6.25e-6 1,000 225,000
Whisper-small 1 4 1.25e-5 1,000 225,000
Table 19: Hyperparameters for training speaker verification models.
Model Batch size LR Schedule Init LR Base LR Epochs
ECAPA-TDNN 128 Cyclic 5e-3 1e-8 20
ResNet-TDNN 128 Cyclic 5e-3 1e-8 20
x-vector 128 Linear 5e-3 1e-4 20
Table 20: Hyperparameters for training speaker editing models.
Model Params Batch size LR Warmup Training steps
CampNet 21.22M 16 2e-4 8,000 200,000
EditSpeech 48.15M 16 2e-4 8,000 200,000
A3T 14.86M 16 2e-4 8,000 200,000
FluentSpeech 23.86M 30 2e-4 8,000 200,000